OpenStudy (anonymous):

A quadrilateral contains two equal sides measuring 12 cm each and an included right angles. If the measure of the third side is 8 cm and the angle opposite the right angle is 120, find the measure of the fourth side and the area of the quadrilateral.

OpenStudy (mathstudent55):



OpenStudy (ranga):



OpenStudy (ranga):

You can find AC from the right triangle ADC. Assume BC = x. You can use the law of cosines to to find AC in terms of x. Equate the above two equations and solve for x.

OpenStudy (ranga):

AC^2 = AD^2 + DC^2 = 12^2 + 12^2 = 288 -------- (1) AC^2 = AB^2 + x^2 - 2AB(x)cos(120) = 8^2 + x^2 - (2)(8)(x)cos(120) = 64 + x^2 + 8x AC^2 = x^2 + 8x + 64 ----------- (2) Equate (1) and (2) and solve for x

Directrix (directrix):

From last night: http://openstudy.com/users/directrix#/updates/5275f0c5e4b0f6d98a1f7ae9